1. We are arranging a series of components within the cockpit of an airplane. To avoid congestion of controls, some switches have been placed on an additional panel placed overhead of the seated pilots. The average functional, seated overhead reach for males is 1.264 m with a standard deviation of 0.085 m. Females have a reach of 1.198 m with a standard deviation of 0.066 m. Our latest design iteration requires a functional reach demand of 1.29 m. For both male and female population draw the distribution curve and indicate the value and placement of the population mean and task demand level. What percentage of males and females are capable of performing this task? Is this a good design? Explain. What design strategy will you adopt and why? a. b. c.
Answer #1

a.

2-12 m Task recent of male 09Y79 lv

b.

Percentage of male capable performing the task = P(Male reach > 1.29 meter)

= P[Z > (1.29 - 1.264) / 0.085] = P[Z > 0.3059] = 0.3798 = 37.98%

Percentage of female capable performing the task = P(Female reach > 1.29 meter)

= P[Z > (1.29 - 1.198) / 0.066] = P[Z > 1.394] = 0.0817 = 8.17%

c.

Since the percentage of male/female capable of performing the task is low, this is not a good design.

For design in which at least 95% of male/female are capable of performing the task, the task demand level should be at

1.198 - 1.96 * 0.066 = 1.06864 meter (z value for 95% confidence is 1.96)